How many pounds of 80-cent nuts and how many pounds of 50-cent nuts must a dealer use to produce a mixture of 90-pounds to sell at 75-cents per pound?
Answer by Alan3354(69443)

Do it like this one:
----------
how many liters of a 70% acid solution must be mixed with a 15% acid solution to get 385L of a 60% acid solution?
-------------------
s = amount of 70%
f = amount of 15%
-------
s + f = 385 (total solution)
0.7s + 0.15f = 0.6*385 (total acid)
-------
f = 385 - s
Sub for f in 2nd equation
0.7s + 0.15(385-s) = 231
70s + 5775 - 15s = 23100
55s = 17325
s = 315 liters of 70%
------------
f = 70